OpenStudy (rissaunnie):

urgent FAN AND MEDAL a country's population in 1991 was 147 million. In 1998 it was 153 million what's the population in 2017

OpenStudy (anonymous):

are you supposed to use exponential growth as in \[P(t)=P_0e^{rt}\]?

OpenStudy (anonymous):

or can you do something easier?

OpenStudy (rissaunnie):

its like P=Ae^kt

OpenStudy (anonymous):

we can do it that way if you like take a bit of time

OpenStudy (rissaunnie):

sure

OpenStudy (anonymous):

we can also do it a quick way using just the numbers given lets to it that way first

OpenStudy (rissaunnie):

alright

OpenStudy (anonymous):

in 7 year (from 1991 to 1998) it went from 147 to 153 you can model it as \[P(t)=147\times \left(\frac{153}{147}\right)^{\frac{t}{7}}\]

OpenStudy (anonymous):

then if you want to know what it is 26 years later, from 1991 to 2017 replace \(t\) by \(26\) and get this http://www.wolframalpha.com/input/?i=147%28153%2F147%29^%2826%2F7%29

OpenStudy (anonymous):

would you like to do it the slow way now?

OpenStudy (anonymous):

set it up as \[P=Ae^{kt}\] and find \(k\) ?

OpenStudy (anonymous):

you might have to do it on a test or sommat, i don't know it is up to you

OpenStudy (rissaunnie):

since its timed i might use wolfram i just dont know how to enter the problem into it

OpenStudy (rissaunnie):

thats y i didnt use it before

OpenStudy (anonymous):

you can use the link i sent

OpenStudy (anonymous):

otherwise we have to set \[153=147e^{9k}\] and solve for \(k\) then make the equation and evaluate at \(t=26\)

OpenStudy (heyitszia1):

Substitute the original population and the model becomes \[P = 147e^{kt}\] 7 years later the population becomes 153. Substitute P with 153. \[153 = 147e^{7k}\] Now solve for k. \[153 = 147e^{7k}\] and t = 16

OpenStudy (anonymous):

that is the second way do you know how to solve \[153=147e^{7k}\] for \(k\)?

OpenStudy (rissaunnie):

no

OpenStudy (anonymous):

do you want to know or are you satisfied with the answer above?